OpenStudy (anonymous):

If you were asked to solve this system of equations by adding, which of the following would be the best first step? 2x + y = 13 3x + y = 18

OpenStudy (angela210793):

subtract both sides of the eq 2x+y-3x-y=13-18 -x=-5 x=5 y=13-10=3